The City of Cambridge is concerned about the number of wild MIT students who will be partying on Halloween and wants to limit the number of parties in order to curb the costs of policing underage drinking and noise violations. The city has asked you to evaluate the welfare implications of the policies they are considering. Think of price here as the amount of money party hosts will collect at the door from party-goers. The demand for parties is given by ?=300−2?. The supply of parties is given by ?=?.

Find the equilibrium price and quantity in the market for Halloween parties.

Now suppose that the city has found a fail-safe way to tax parties by requiring each party host to pay a fee of $60. Find the new price for party-goers, the after-tax price received by the party hosts, and the new equilibrium quantity. (Do not round.)

Let ?? be the price the suppliers receive (after the tax), ?? be the price the consumers pay for each party, and ?? be the after-tax quantity.
Answer #1

At equilibrium, demand = supply,
300−2? = p
p = 100
Q = 100

new supply curve after tax, P = Q+60
At equilibrium, demand = supply,
300−2? = p-60
p = 120
Q = 60
Pc = 120, Ps = 60, Qt = 60